OpenStudy (anonymous):

f(x)=-x + 2 g(x)=-4x^2+ x + 5 Find (fog)(x)

OpenStudy (anonymous):

@jim_thompson5910 whats fog?

OpenStudy (cwrw238):

same as fg(x)

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so first i'd do -4(-x+2)^2+(-x+2)+5

OpenStudy (cwrw238):

substiitute g(x) for the x in f(x) and simplify

OpenStudy (anonymous):

ok so my final answer would be -4x^2+15x-9

OpenStudy (anonymous):

actually i would do -(-4x^2+x+5)+2

OpenStudy (anonymous):

and my final answer would be 4x^2-x-3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Can you please check this? last one! @kirbykirby @jim_thompson5910

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

someone hook a girl up?!?!?!?!

OpenStudy (kirbykirby):

yes the last answer you gave is good

jimthompson5910 (jim_thompson5910):

( f o g )(x) = f(g(x)) f(x) = -x + 2 f(g(x)) = -(g(x)) + 2 f(g(x)) = -(-4x^2+ x + 5) + 2 simplify

OpenStudy (anonymous):

thank you all!!!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

so you got it, the final answer is 4x^2-x-3 (just saw you posted that), so nice work

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!!!